OpenStudy (anonymous):

Two dice are thrown. If each face is equally likely to turn up, find the probability that the sum is at least 7. (Write your answer as a fraction.)

OpenStudy (anonymous):

at least 7 means 7, 8, 9, 10, 11 or 12

OpenStudy (anonymous):

take the probabilities and add them up keep 36 as your denominator typo should be \[\frac{6}{36}+\frac{5}{36}+\frac{4}{36}+\frac{3}{36}+\frac{2}{36}+\frac{1}{36}\]

OpenStudy (anonymous):

thank you